# Natural Language Processing for Dialects of a Language: A Survey

### v1

@@ 2.2. Rethinking LLM benchmarks
Yet, the corpora used to train language models and more importantly, the datasets used to evaluate them do not necessarily reflect this gap. In the case of GPT-4, the evaluation dataset consists of questions from the MMLU benchmark written in Standard American English [^8] GPT-4 also claims multilingual performance in various languages using translations of English questions into target languages. Using translations of sentences to claim multilingual ability is inaccurate (Joshi et al., 2023). Standard benchmarks used to claim performance of a language model for English primarily contain Standard American English. It has been found that the performance does not extend to NLU tasks for dialects of English (Ziems et al., 2022). This holds for most foundation models that are trained on large amounts of data. The distribution of languages in the training corpora is either not known or difficult to determine. There are variants of English, i.e., dialects, that these models brush under the carpet, despite the existence of corpora of regional varieties of English [^9] https://www.ice-corpora.uzh.ch/en.html; Accessed on 9th January, 2024. and of corpora of English as a lingua franca [^10] https://www.kielipankki.fi/corpora/elfa/; Accessed on 9th January, 2024.. Some examples showing the impact of dialects on the performance of NLP tasks are presented in Table 2.3. We note that these papers are from the past few years, which have otherwise witnessed a great development in the reported performance of NLP models.
@@ 2.3. Fair and equitable technologies
NLP Task Paper Impact
@@
classification
 (Blodgett et al., 2016) Language detection shows lower performance
@@
classification
 (Okpala et al., 2022) Text in African-American English may be
predicted higher as hate speech.
@@
Understanding
 (Ziems et al., 2022) Popular models perform worse on GLUE tasks

### v2

@@ Abstract.
Blue text macro. Red text macro. Black text macro.
@@ 2.2. Rethinking LLM benchmarks
Yet, the corpora used to train language models and more importantly, the datasets used to evaluate them do not necessarily reflect this gap. (Inoue et al., 2021) examine the performance of BERT-based models for varieties/dialects of Arabic, and show that dialect proximity of pre-training and fine-tuning data bears impact on the performance of the downstream task. In the case of GPT-4, the evaluation dataset consists of questions from the MMLU benchmark written in Standard American English [^8] GPT-4 also claims multilingual performance in various languages using translations of English questions into target languages. Using translations of sentences to claim multilingual ability is inaccurate (Joshi et al., 2023). Standard benchmarks used to claim performance of a language model for English primarily contain Standard American English. It has been found that the performance does not extend to NLU tasks for dialects of English (Ziems et al., 2022). This holds for most foundation models that are trained on large amounts of data. The distribution of languages in the training corpora is either not known or difficult to determine. There are variants of English, i.e., dialects, that these models brush under the carpet, despite the existence of corpora of regional varieties of English [^9] https://www.ice-corpora.uzh.ch/en.html; Accessed on 9th January, 2024. and of corpora of English as a lingua franca [^10] https://www.kielipankki.fi/corpora/elfa/; Accessed on 9th January, 2024.. Some examples showing the impact of dialects on the performance of NLP tasks are presented in Table 1. We note that these papers are from the past few years, which have otherwise witnessed a great development in the reported performance of NLP models.
@@ 2.3. Fair and equitable technologies
  4. (4)
    Prejudice in prediction of employability and criminality: (Hofmann et al., 2024) show that dialects may introduce bias in the output of language models. As a result, a person’s output with respect to their employability or criminality may be affected based on the dialects they use.
@@
{tblr}
width = colspec = Q[225]Q[257]Q[617], hline1,8 = -0.08em, hline2,5 = -0.05em, NLP Task & Paper Impact
@@
classification  (Blodgett et al., 2016) Language detection shows lower performance
@@
classification  (Okpala et al., 2022) Text in African-American English may be
predicted more commonly as hate speech.
@@
Understanding  (Ziems et al., 2022) Popular models perform worse on GLUE tasks